# Pre-training Cross-lingual Open Domain Question Answering with Large-scale Synthetic Supervision

### v1

@@ A.3.1 Cross-lingual Passage Retrieval
  • DPR+MT.  This is a translate-test baseline that involves the translation of queries into English during test time, followed by monolingual passage retrieval using the English DPR encoder.
  • CORA.  This method trains a multilingual DPR encoder iteratively, with positive and negative passages identified by a multilingual QA model.
  • Sentri.  An iterative self-training method that uses the latest retriever to identify positive and negative passages through answer string matching for updating the training dataset. It also employs data augmentation techniques by translating English datasets.
  • QuiCK.  A knowledge distillation method that trains a multilingual bi-encoder retriever, learning from a query generator as the teacher. The query generator is also used for generating synthetic multilingual queries to enhance the knowledge distillation process.
  • DrDecr.  A multilingual ColBERT model that learns from an English ColBERT on parallel queries, sourced from both parallel corpora and human-translated gold queries released by XOR-Retrieve.
  • LAPCA.  A pre-training method that takes the first paragraphs of parallel Wikipedia pages as the parallel corpus for cross-lingual pre-training. Additionally, machine translation is employed for data expansion.
@@ A.3.2 Multilingual Open Domain Question Answering
  • mDPR.  This model uses a multilingual reader to identify the answer span from the passages retrieved by the mDPR retriever as described above.
  • MT+DPR.  This represents the translate-test baseline, in which queries are translated into English and the answers are identified within English passages retrieved by the DPR+MT retriever. The English answer is then translated back to the target language if necessary.
  • ReAtt+MT.  This is the English teacher employed in our stage-1 pre-training. We use a state-of-the-art machine translation model [^12] https://huggingface.co/facebook/m2m100_418M to translate the queries into English at test time. It always retrieves passages from English Wikipedia and generates answers in English. The generated answer is translated back to the target language if necessary.
  • GMT+GS.  This pipeline follows the same procedure as MT+DPR except that we employ Google Search for passage retrieval and Google Machine Translation services for query and answer translation.
  • Monolingual baseline (BM25).  Instead of using a multilingual DPR or an English DPR model with query translation, this baseline always retrieves the passage from the target language and extracts the answer using a multilingual reader.
  • MT+Mono.  This is a combination of the BM25 and MT+DPR baselines, which first does monolingual QA for the target language using the BM25 method and resorts to the MT+DPR baseline if no answer is found.
  • Fusion-in-Decoder.  This encompasses a family of multilingual retrieval-augmented generation models, which take the passages returned by a multilingual retriever as inputs to generate the answer in the target language. CORA, Sentri and LAPCA are included in this family by using the passages returned by their respective retrievers.
Приложение B Detailed Zero-shot Evaluation
Cross-lingual Retrieval
Table 4 presents the detailed result comparisons in each of the 20 unseen languages covered by MKQA. Notably, CLASS-ZS outperforms other baselines significantly on average and achieves the best results in nearly all languages except for Vietnamese. Comparing the three variants of our method, fine-tuning on supervised English data significantly enhances cross-lingual transfer abilities to every unseen language (i.e., CLASS-US vs CLASS-ZS). However, fine-tuning CLASS-ZS on a limited number of supervised multilingual data with a restricted language set does not lead to improved generalization performance, as indicated by the result comparison in every language between CLASS-ZS and CLASS. Furthermore, a decrease in performance is also observed in both supervised and zero-shot settings when either stage-2 pre-training or the entire pre-training procedures are omitted, highlighting the effectiveness of our pre-training approach in enhancing cross-lingual ability.
Multilingual QA
Table 5 presents the detailed multilingual QA results for each of the 20 unseen languages covered by MKQA. We observe similar patterns where CLASS-US surpasses a range of machine-translation-based methods and CLASS-ZS outperforms the supervised CORA with a significant margin. Further fine-tuning CLASS-ZS on a limited number of supervised multilingual data with a restricted language set hampers its generalizability, with a decline in performance across all examined languages.
Monolingual Retrieval
We evaluate the model on the monolingual retrieval setting, with the results on the Mr. TyDi dataset (Zhang et al., 2021) reported in Table 6. We compare CLASS-ZS with a series of zero-shot multilingual retrievers that have been fine-tuned on English datasets. Notably, even though CLASS-ZS have not undergone explicit training for monolingual retrieval, it consistently outperforms the other models in every language examined, with the performance gains being particularly pronounced in languages written in Non-Latin scripts (e.g., Japanese, Korean, and Thai).
@@
  - Stage-2 56.9 57.3 57.2 57.0 47.3 51.8 56.2 42.9 57.6 58.7 56.0 55.3 55.5 56.8 56.1 53.3 51.5 51.4 49.9 49.4 53.9
  - Pre-train 56.5 55.3 55.9 55.1 44.8 50.8 55.0 41.3 56.4 57.4 55.8 53.3 54.8 56.5 53.7 51.9 49.6 47.3 46.4 45.8 52.2
Таблица 4: Zero-shot cross-lingual retrieval results (R@2kt) on the MKQA dataset. "cn": "Zh-cn"(Chinese, simplified). "hk": "Zh-hk"(Chinese, Hong Kong). "tw": "Zh-tw"(Chinese, traditional).
@@
Таблица 5: Zero-shot multilingual question answering results (F1) on the MKQA dataset. "cn": "Zh-cn"(Chinese, simplified). "hk": "Zh-hk"(Chinese, Hong Kong). "tw": "Zh-tw"(Chinese, traditional).
Method Ar Bn Fi Id Ja Ko Ru Sw Te Th Avg
mDPR 62.0 67.1 37.5 46.6 53.5 49.0 49.8 26.4 35.2 45.5 47.3
LAPCA 89.4 95.0 83.6 90.6 78.1 76.0 80.2 74.4 93.0 92.8 85.3
mContriever 88.7 91.4 88.1 89.8 81.7 78.2 83.8 91.4 96.6 90.5 88.0
CLASS-ZS 93.8 96.4 92.6 94.8 91.3 86.8 92.5 93.1 98.8 97.1 93.7
Таблица 6: Zero-shot monolingual retrieval with results (R@100) on the test set of Mr. TyDi dataset.
Приложение C More Analysis
Рис. 6: Performance evolution in stage-1 pre-training.
Performance Evolution during Pre-training
Рис. 7: Scaling training data on cross-lingual retrieval.
Few-Shot Cross-lingual Retrieval
Рис. 8: Effects of employing different numbers of retrieved passages for QA during inference time.
Effects of Number of Retrieved Passages
You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "Strapping Young Lad (lyh. SYL) oli Devin Townsendin vuonna 1994 perustama kanadalainen metalliyhtye."into a natural question whose question word is "Milloin"and answer is "1994". Please respond in the format: "The transformed question is: Milloin Devin Townsend perusti kanadalaisen metalliyhtyeen Strapping Young Lad (lyh. SYL)?"
You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence " В 215 году Цао Цао атаковал Чжан Лу и разгромил его в битве в проходе Янпингуань. "into a natural question whose question word is " Кто "and answer is " Чжан Лу ". Please respond in the format: "The transformed question is: Кто был атакован Цао Цао и разгромлен в битве в проходе Янпингуань в 215 году? "
You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "熊野那智神社(くまのなちじんじゃ)は、宮城県名取市にある神社である。"into a natural question whose question word is "どこ"and answer is "宮城県". Please respond in the format: "The transformed question is: 熊野那智神社はどこにある神社ですか?"
You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "19세기 후반에 아일랜드에는 독립과 토지개혁을 요구하는 운동이 크게 확산되었다."into a natural question whose question word is "어디"and answer is "아일랜드". Please respond in the format: "The transformed question is: 19세기 후반에 독립과 토지개혁을 요구하는 운동이 크게 확산된 나라는 어디입니까?"
You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence " ولكن في مجال التعليم العالي تفتقر المسلمين، ونظرا لأفراد أسرهم وأقاربهم أخذها عن وظائف في دول الخليج وجنوب شرق آسيا (أساسا وسنغافورة، وألمانيا ماليزيا وبروناي في جنوب آسيا) نفسها في سن مبكرة. "into a natural question whose question word is " أين "and answer is " جنوب شرق آسيا ". Please respond in the format: "The transformed question is: أين يأخذ أفراد أسر المسلمين وأقاربهم وظائف غالبًا، مما يؤدي إلى فقدان الاهتمام بالتعليم العالي؟ "
You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "bharte Hajar Hajar manuSh AnaHaer mara JaJ, ikn/tu dhr/mpRcarkra taedr pRit Udasiin."into a natural question whose question word is "ekathaJ"and answer is "bhart". Please respond in the format: "The transformed question is: ekathaJ Hajar Hajar manuSh AnaHaer mara JaJ EbNNG dhr/mpRcarkra taedr pRit Udasiin thaek?"
Рис. 9: Meta-examples obtained by prompting ChatGPT are shown for each language coverd by Xor-TyDi QA. Lightblue texts indicate the transformed questions.
Приложение D Query Transformation Examples
Table 9 showcases examples illustrating the generation of meta-examples through prompting ChatGPT. Prompts D, D, D, D, D, D, and D provide detailed illustrations of prompting a much smaller large language model, LLaMA-2-7B, to perform query transformation using In-Context Learning, which incorporates meta-examples into the prompt to guide the model’s behaviour. The choice of the question word is determined based on the detected entity type of the answer and the heuristic rules outlined in Table 7.
@@
Таблица 7: The heuristics rules for choosing the most appropriate question word based on named entity types (taken from Lewis et al. (2019)).
@@
Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Toisaalta hän oli taiteiden suosija ja hänen valtakaudellaan Preussi sai haltuunsa suuren osan Puola-Liettuasta Puolan jaoissa vuosina 1793 ja 1795. Question word: Missä Answer: Preussi Transformed Question: Missä maassa taiteiden suosija hallitsi ja missä valtakunnassa saatiin haltuunsa suuri osa Puola-Liettuasta Puolan jaoissa vuosina 1793 ja 1795? Sentence: Hän pelasi urallaan myös Ruotsissa ja Slovakiassa. Question word: Missä Answer: Slovakia Transformed Question: Missä maassa hän pelasi urallaan Ruotsin lisäksi? Sentence: Barokin jälkeen concerto grossoja ovat säveltäneet muun muassa Heitor Villa-Lobos, Bohuslav Martinů, Alfred Schnittke ja Philip Glass. Question word: Kuka Answer: Bohuslav Martinů Transformed Question: Kuka säveltäjistä Heitor Villa-Lobosin, Alfred Schnittken ja Philip Glassin ohella on säveltänyt concerto grossoja barokin jälkeen? Sentence: Hänen ajatteluunsa vaikuttivat muun muassa buddhalaiset ja taolaiset ideat, joihin hän tutustui Aasian matkoillaan, Mahatma Gandhin väkivallattomuusliike, sekä hänen katolinen uskontonsa. Question word: Kuka Answer: Mahatma Gandhi Transformed Question: Kuka vaikutti hänen ajatteluunsa, mahtimaailmaan ja katoliseen uskontonsa? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: On the other hand, he/she was a fan of the arts and during his/her reign, Prussia took over a large part of Poland-Lithuania in the partitions of Poland in 1793 and 1795. Question word: Where Answer: Prussia Transformed Question: In which country did the lover of the arts rule and in which kingdom was a large part of Poland-Lithuania taken over during the partitions of Poland in 1793 and 1795? Sentence: He/She also played in Sweden and Slovakia during her career. Question word: Where Answer: Slovakia Transformed Question: In which country did he/she play in his/her career besides Sweden? Sentence: After the Baroque, concerto grossos have been composed by, among others, Heitor Villa-Lobos, Bohuslav Martinů, Alfred Schnittke and Philip Glass. Question word: Kuka Answer: Bohuslav Martinů Transformed Question: Besides Heitor Villa-Lobos, Alfred Schnittke and Philip Glass, which of the composers has composed concerto grossos after the Baroque? Sentence: His/Her thinking was influenced, among other things, by Buddhist and Taoist ideas, which he/she got to know during his/her travels in Asia, Mahatma Gandhi’s non-violence movement, and his/her Catholic religion. Question word: Who Answer: Mahatma Gandhi Transformed Question: Who influenced his/her thinking, the world of power and his/her Catholic religion?
@@
Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Корабли проекта выполняли контроль за учениями ВМС стран НАТО в Норвежском и Средиземном морях, следили за корабельными и авианосными группами флотов США и Великобритании. Question word: Кто Answer: НАТО Transformed Question: Кто выполнял контроль за учениями ВМС в Норвежском и Средиземном морях и следил за корабельными и авианосными группами флотов США и Великобритании? Sentence: 1 апреля 1768 года Доверню назначают пенсию Королевской академии музыки в размере 1000 ливров как автору музыки. Question word: Кто Answer: Королевской академии музыки Transformed Question: Кто 1 апреля 1768 года назначил пенсию в размере 1000 ливров Доверню как автору музыки? Sentence: Софи́я Шарло́тта Авгу́ста (22 февраля 1847, Мюнхен — 4 мая 1897, Париж) — принцесса Баварская, герцогиня Баварская, позднее герцогиня Алансонская и Орлеанская. Question word: Где Answer: Мюнхен Transformed Question: Где родилась София Шарлотта Августа, принцесса Баварская? Sentence: В первой половине XIX века паровозы в Россию, в основном, ввозились из-за рубежа. Question word: Когда Answer: XIX век Transformed Question: Когда паровозы в Россию, в основном, ввозились из-за рубежа? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: The project’s ships monitored NATO naval exercises in the Norwegian and Mediterranean Seas and monitored ship and aircraft carrier groups of the US and British navies. Question word: Who Answer: NATO Transformed Question: Who monitored naval exercises in the Norwegian and Mediterranean seas and monitored ship and aircraft carrier groups of the US and British fleets? Sentence: On April 1, 1768, Dauvergne was awarded a pension from the Royal Academy of Music in the amount of 1000 livres as the author of music. Question word: Who Answer: Royal Academy of Music Transformed Question: Who, on April 1, 1768, awarded a pension of 1000 livres to Dovergne as the author of music? Sentence: Sophia Charlotte Auguste (22 February 1847, Munich - 4 May 1897, Paris) - Princess of Bavaria, Duchess of Bavaria, later Duchess of Alençon and Orléans. Question word: Where Answer: Munich Transformed Question: Where was Sophia Charlotte Augusta, Princess of Bavaria born? Sentence: In the first half of the 19th century, steam locomotives were mainly imported to Russia from abroad. Question word: When Answer: 19th century Transformed Question: When were steam locomotives mainly imported into Russia from abroad?

### v3

@@ A.3.1 Cross-lingual Passage Retrieval
  • DPR+MT  (Asai et al., 2021a). This is a translate-test baseline that involves the translation of queries into English during test time, followed by monolingual passage retrieval using the English DPR encoder.
  • CORA  (Asai et al., 2021b). This method trains a multilingual DPR encoder iteratively, with positive and negative passages identified by a multilingual QA model.
  • Sentri  (Sorokin et al., 2022). An iterative self-training method that uses the latest retriever to identify positive and negative passages through answer string matching for updating the training dataset. Machine translation is used for data augmentation.
  • QuiCK  (Ren et al., 2022). A knowledge distillation method that trains a multilingual bi-encoder retriever, learning from a query generator as the teacher. The query generator is also used for generating synthetic multilingual queries to enhance knowledge distillation.
  • DrDecr  (Li et al., 2022). A multilingual ColBERT model that learns from an English ColBERT on parallel queries, sourced from both parallel corpora and human-translated gold queries released by XOR-Retrieve.
  • LAPCA  (Abulkhanov et al., 2023). A pre-training method that takes the first paragraphs of parallel Wikipedia pages as the parallel corpus for cross-lingual pre-training, with augmented data through machine translation.
  • SWIM-X  (Thakur et al., 2023). A method that uses large language models to generate synthetic queries from unlabelled corpus with textual summary generation as an intermediate step. A multilingual dense retrieval model is fine-tuned exclusively on synthetic data.
@@ A.3.2 Multilingual Open Domain Question Answering
  • MT+DPR  (Asai et al., 2021a). This represents the translate-test baseline, in which queries are translated into English and the answers are identified within English passages retrieved by the DPR+MT retriever. The English answer is then translated back to the target language if necessary.
  • ReAtt+MT  (Jiang et al., 2022). This is the English teacher employed in the cross-lingual retrieval pre-training. We use a state-of-the-art machine translation model [^11] https://huggingface.co/facebook/m2m100_418M to translate the queries into English at test time. It always retrieves passages from English Wikipedia and generates answers in English. The generated answer is translated back to the target language.
  • GMT+GS  (Asai et al., 2021a). This pipeline follows the same procedure as MT+DPR except that we employ Google Search for passage retrieval and Google Machine Translation services for query and answer translation.
  • Monolingual baseline (BM25)  (Asai et al., 2021a). Instead of using a multilingual DPR or an English DPR model with query translation, this baseline always retrieves the passage from the target language and extracts the answer using a multilingual reader.
  • MT+Mono  (Asai et al., 2021a). This is a combination of the BM25 and MT+DPR baselines, which first does monolingual QA for the target language using the BM25 method and resorts to the MT+DPR baseline if no answer is found.
  • Fusion-in-Decoder.  This encompasses a family of multilingual retrieval-augmented generation models, which take the passages returned by a multilingual retriever as inputs to generate the answer in the target language. CORA (Asai et al., 2021b), Sentri (Ren et al., 2022) and LAPCA (Abulkhanov et al., 2023) are included in this family by using the passages returned by their respective retrievers.
@@
- MLQA-PT 56.9 57.3 57.2 57.0 47.3 51.8 56.2 42.9 57.6 58.7 56.0 55.3 55.5 56.8 56.1 53.3 51.5 51.4 49.9 49.4 53.9
- Pre-train 56.5 55.3 55.9 55.1 44.8 50.8 55.0 41.3 56.4 57.4 55.8 53.3 54.8 56.5 53.7 51.9 49.6 47.3 46.4 45.8 52.2
Таблица 5: Zero-shot cross-lingual retrieval results (R@2kt) on the MKQA dataset. "cn": "Zh-cn"(Chinese, simplified). "hk": "Zh-hk"(Chinese, Hong Kong). "tw": "Zh-tw"(Chinese, traditional).
@@
Таблица 6: Zero-shot multilingual question answering results (F1) on the MKQA dataset. "cn": "Zh-cn"(Chinese, simplified). "hk": "Zh-hk"(Chinese, Hong Kong). "tw": "Zh-tw"(Chinese, traditional).
Приложение B Detailed Zero-shot Evaluation
Cross-lingual Retrieval.
Table 5 presents the detailed result comparisons in each of the 20 unseen languages covered by MKQA. Notably, CLASS-ZS outperforms other baselines significantly on average and achieves the best results in nearly all languages except for Vietnamese. Comparing the three variants of our method, fine-tuning on supervised English data significantly enhances cross-lingual transfer abilities to every unseen language (i.e., CLASS-US vs CLASS-ZS). However, fine-tuning CLASS-ZS on a limited number of supervised multilingual data with a restricted language set does not lead to improved generalization performance, as indicated by the result comparison in every language between CLASS-ZS and CLASS. Furthermore, a decrease in performance is also observed in both supervised and zero-shot settings when either multilingual QA pre-training or the entire pre-training procedures are omitted, highlighting the effectiveness of our pre-training approach in enhancing cross-lingual ability.
Multilingual QA.
Table 6 presents the detailed multilingual QA results for each of the 20 unseen languages covered by MKQA. We observe similar patterns where CLASS-US surpasses a range of machine-translation-based methods and CLASS-ZS outperforms the supervised CORA by a significant margin. Further fine-tuning CLASS-ZS on a limited number of supervised multilingual data with a restricted language set hampers its generalizability, with a decline in performance across all examined languages.
Model F1 EM BLEU
CLASS 30.4 21.0 20.6
CLASS w/o MLQA-PT 30.2 21.4 20.3
CLASS w/o Pre-train 29.7 20.9 19.8
Таблица 7: Multilingual QA results on queries requiring cross-lingual evidence retrieval.
Приложение C Error Analysis
We add additional error analysis regarding the issue identified in multilingual QA (i.e., XOR-Full).
Model R@2kt R@5kt R@10kt
CLASS 61.0 70.6 75.6
CLASS w/o MLQA-PT 59.0 68.8 74.6
CLASS w/o Pre-train 50.6 59.3 65.4
(a) English or target language answer is in top-n retrieved tokens.
Model R@2kt R@5kt R@10kt
CLASS 50.4 63.1 70.8
CLASS w/o MLQA-PT 46.3 59.2 67.8
CLASS w/o Pre-train 32.7 42.1 50.4
(b) Only English answer is in top-n retrieved tokens.
Model R@2kt R@5kt R@10kt
CLASS 41.8 47.3 50.8
CLASS w/o MLQA-PT 42.7 47.9 51.2
CLASS w/o Pre-train 41.0 46.9 50.4
(c) Only Target language answer is in top-n retrieved tokens.
Таблица 8: Retrieval accuracy of queries requiring answers based on English evidence.
Contain English Ans No English Ans
Contain Target Ans F1: 41.0/EM: 30.6/BLEU: 33.2 F1: 37.7/EM: 28.3/BLEU: 32.5
No Target Ans F1: 13.5/EM: 2.1/BLEU: 12.9 F1: 10.1/EM: 1.0/BLEU: 6.8
Таблица 9: Multilingual QA results on queries requiring cross-lingual evidence retrieval, grouped by whether the gold-standard answer string in English or the target language appears within the top-n retrieved tokens.
C.1 Quantitative Analysis
Our focus is on analysing the behaviour of our model when handling cross-lingual queries in XOR-Full. These queries require answers based on English evidence (Asai et al., 2021a). Initially, we analyse the retrieval accuracy of our model by assessing whether the top-n retrieved tokens contain the answer string in English or the target language.
As shown in Table 8, our pre-training method shows significant improvements in finding correct English evidence for those queries requiring cross-lingual evidence retrieval (e.g., 50.4% -> 70.8%) while maintaining competitive performance (Table 8(c)) in finding in-language (i.e., the question language) evidence if there exists. Nevertheless, we have observed that these advancements do not translate into enhancements in the subsequent QA task, wherein the model is supposed to produce an answer in the same language as the question with English supporting documents. Table 7 shows that our complete CLASS model fails to achieve additional benefits in QA tasks despite its outstanding performance in retrieving cross-lingual evidence.
To gain deeper insights into the behaviour of our model, we specifically analyse its QA performance whenever the top-n retrieved evidence contains the gold answer in either English or the target language. As indicated in Table 9, our model demonstrates reasonable performance only when the correct answer string is presented in the target language. However, it often fails to generate the correct answer when the gold standard answer is provided solely in English, despite our model being able to include the correct English answer in its top-10k retrieved tokens 71% of the time. This indicates a deficiency in our model’s ability to identify correct clues for QA among cross-lingual evidence. An example is shown in Figure LABEL:fig:ca_score. In cases where the top 100 retrieved passages contain answer strings in the target language, our model tends to assign significantly higher scores to passages containing these target language answer strings. By contrast, when only English answer strings are present, the distribution of cross-attention scores across all retrieved passages becomes more uniform, leading to a general narrowing of the gap between positively relevant passages and irrelevant ones.
Таблица 10: An example of our model in finding correct evidence while failing to generate the right answer in the target language.
@@ C.1 Quantitative Analysis
Таблица 11: The heuristics rules for choosing the most appropriate question word based on named entity types (taken from Lewis et al. (2019)).
C.2 Case Study
As shown in Table 10, our model successfully retrieves the appropriate supporting document as its top-1 retrieval. However, it encounters challenges in generating Telugu answers, whereas it performs accurately in English. This highlights our model’s inability to translate English evidence into answers in the target language, necessitating further efforts to enhance the model’s capabilities in cross-lingual evidence reasoning and answer generation.
Приложение D More Analysis
Рис. 8: Performance evolution in stage-1 pre-training.
Performance Evolution during Pre-training.
Рис. 9: Scaling training data on cross-lingual retrieval.
Few-Shot Cross-lingual Retrieval.
Рис. 10: Effects of employing different numbers of retrieved passages for QA during inference time.
Effects of Number of Retrieved Passages.
Finnish Prompt You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "Strapping Young Lad (lyh. SYL) oli Devin Townsendin vuonna 1994 perustama kanadalainen metalliyhtye."into a natural question whose question word is "Milloin"and answer is "1994". Please respond in the format: "The transformed question is: Milloin Devin Townsend perusti kanadalaisen metalliyhtyeen Strapping Young Lad (lyh. SYL)?"
Russian Prompt You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence " В 215 году Цао Цао атаковал Чжан Лу и разгромил его в битве в проходе Янпингуань. "into a natural question whose question word is " Кто "and answer is " Чжан Лу ". Please respond in the format: "The transformed question is: Кто был атакован Цао Цао и разгромлен в битве в проходе Янпингуань в 215 году? "
Japanese Prompt You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "{CJK}UTF8min熊野那智神社(くまのなちじんじゃ)は、宮城県名取市にある神社である。"into a natural question whose question word is "{CJK}UTF8minどこ"and answer is "{CJK}UTF8min宮城県". Please respond in the format: "The transformed question is: {CJK}UTF8min熊野那智神社はどこにある神社ですか?"
Korean Prompt You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "{CJK}UTF8mj19세기 후반에 아일랜드에는 독립과 토지개혁을 요구하는 운동이 크게 확산되었다."into a natural question whose question word is "{CJK}UTF8mj어디"and answer is "{CJK}UTF8mj아일랜드". Please respond in the format: "The transformed question is: {CJK}UTF8mj19세기 후반에 독립과 토지개혁을 요구하는 운동이 크게 확산된 나라는 어디입니까?"
Arabic Prompt You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence " ولكن في مجال التعليم العالي تفتقر المسلمين، ونظرا لأفراد أسرهم وأقاربهم أخذها عن وظائف في دول الخليج وجنوب شرق آسيا (أساسا وسنغافورة، وألمانيا ماليزيا وبروناي في جنوب آسيا) نفسها في سن مبكرة. "into a natural question whose question word is " أين "and answer is " جنوب شرق آسيا ". Please respond in the format: "The transformed question is: أين يأخذ أفراد أسر المسلمين وأقاربهم وظائف غالبًا، مما يؤدي إلى فقدان الاهتمام بالتعليم العالي؟ "
Bengali Prompt You are an AI model that rewrites sentences into questions, using a given question word and answer. Rewrite this sentence "bharte Hajar Hajar manuSh AnaHaer mara JaJ, ikn/tu dhr/mpRcarkra taedr pRit Udasiin."into a natural question whose question word is "ekathaJ"and answer is "bhart". Please respond in the format: "The transformed question is: ekathaJ Hajar Hajar manuSh AnaHaer mara JaJ EbNNG dhr/mpRcarkra taedr pRit Udasiin thaek?"
Telugu Prompt Refer to caption
Рис. 11: Meta-examples obtained by prompting ChatGPT are shown for each language coverd by Xor-TyDi QA. Lightblue texts indicate the transformed questions.
Приложение E Query Transformation Examples
Figure 11 showcases examples illustrating the generation of meta-examples through prompting ChatGPT. Prompts E, E, E, E, E, E, and E provide detailed illustrations of prompting a much smaller large language model, LLaMA-2-7B, to perform query transformation using In-Context Learning, which incorporates meta-examples in the target language $L$ from $\mathbb{K}$ into the prompt to guide the model’s behaviour. The choice of the question word is determined based on the detected entity type of the answer and the heuristic rules outlined in Table 11.
@@ C.2 Case Study
Prompt E.1: Finnish Example & Translation Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Toisaalta hän oli taiteiden suosija ja hänen valtakaudellaan Preussi sai haltuunsa suuren osan Puola-Liettuasta Puolan jaoissa vuosina 1793 ja 1795. Question word: Missä Answer: Preussi Transformed Question: Missä maassa taiteiden suosija hallitsi ja missä valtakunnassa saatiin haltuunsa suuri osa Puola-Liettuasta Puolan jaoissa vuosina 1793 ja 1795? Sentence: Hän pelasi urallaan myös Ruotsissa ja Slovakiassa. Question word: Missä Answer: Slovakia Transformed Question: Missä maassa hän pelasi urallaan Ruotsin lisäksi? Sentence: Barokin jälkeen concerto grossoja ovat säveltäneet muun muassa Heitor Villa-Lobos, Bohuslav Martinů, Alfred Schnittke ja Philip Glass. Question word: Kuka Answer: Bohuslav Martinů Transformed Question: Kuka säveltäjistä Heitor Villa-Lobosin, Alfred Schnittken ja Philip Glassin ohella on säveltänyt concerto grossoja barokin jälkeen? Sentence: Hänen ajatteluunsa vaikuttivat muun muassa buddhalaiset ja taolaiset ideat, joihin hän tutustui Aasian matkoillaan, Mahatma Gandhin väkivallattomuusliike, sekä hänen katolinen uskontonsa. Question word: Kuka Answer: Mahatma Gandhi Transformed Question: Kuka vaikutti hänen ajatteluunsa, mahtimaailmaan ja katoliseen uskontonsa? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: On the other hand, he/she was a fan of the arts and during his/her reign, Prussia took over a large part of Poland-Lithuania in the partitions of Poland in 1793 and 1795. Question word: Where Answer: Prussia Transformed Question: In which country did the lover of the arts rule and in which kingdom was a large part of Poland-Lithuania taken over during the partitions of Poland in 1793 and 1795? Sentence: He/She also played in Sweden and Slovakia during her career. Question word: Where Answer: Slovakia Transformed Question: In which country did he/she play in his/her career besides Sweden? Sentence: After the Baroque, concerto grossos have been composed by, among others, Heitor Villa-Lobos, Bohuslav Martinů, Alfred Schnittke and Philip Glass. Question word: Kuka Answer: Bohuslav Martinů Transformed Question: Besides Heitor Villa-Lobos, Alfred Schnittke and Philip Glass, which of the composers has composed concerto grossos after the Baroque? Sentence: His/Her thinking was influenced, among other things, by Buddhist and Taoist ideas, which he/she got to know during his/her travels in Asia, Mahatma Gandhi’s non-violence movement, and his/her Catholic religion. Question word: Who Answer: Mahatma Gandhi Transformed Question: Who influenced his/her thinking, the world of power and his/her Catholic religion?
@@
Prompt E.2: Russian Example & Translation Rewrite sentences into short and precise questions, using given question words and answers: Sentence: Корабли проекта выполняли контроль за учениями ВМС стран НАТО в Норвежском и Средиземном морях, следили за корабельными и авианосными группами флотов США и Великобритании. Question word: Кто Answer: НАТО Transformed Question: Кто выполнял контроль за учениями ВМС в Норвежском и Средиземном морях и следил за корабельными и авианосными группами флотов США и Великобритании? Sentence: 1 апреля 1768 года Доверню назначают пенсию Королевской академии музыки в размере 1000 ливров как автору музыки. Question word: Кто Answer: Королевской академии музыки Transformed Question: Кто 1 апреля 1768 года назначил пенсию в размере 1000 ливров Доверню как автору музыки? Sentence: Софи́я Шарло́тта Авгу́ста (22 февраля 1847, Мюнхен — 4 мая 1897, Париж) — принцесса Баварская, герцогиня Баварская, позднее герцогиня Алансонская и Орлеанская. Question word: Где Answer: Мюнхен Transformed Question: Где родилась София Шарлотта Августа, принцесса Баварская? Sentence: В первой половине XIX века паровозы в Россию, в основном, ввозились из-за рубежа. Question word: Когда Answer: XIX век Transformed Question: Когда паровозы в Россию, в основном, ввозились из-за рубежа? Rewrite sentences into short and precise questions, using given question words and answers: Sentence: The project’s ships monitored NATO naval exercises in the Norwegian and Mediterranean Seas and monitored ship and aircraft carrier groups of the US and British navies. Question word: Who Answer: NATO Transformed Question: Who monitored naval exercises in the Norwegian and Mediterranean seas and monitored ship and aircraft carrier groups of the US and British fleets? Sentence: On April 1, 1768, Dauvergne was awarded a pension from the Royal Academy of Music in the amount of 1000 livres as the author of music. Question word: Who Answer: Royal Academy of Music Transformed Question: Who, on April 1, 1768, awarded a pension of 1000 livres to Dovergne as the author of music? Sentence: Sophia Charlotte Auguste (22 February 1847, Munich - 4 May 1897, Paris) - Princess of Bavaria, Duchess of Bavaria, later Duchess of Alençon and Orléans. Question word: Where Answer: Munich Transformed Question: Where was Sophia Charlotte Augusta, Princess of Bavaria born? Sentence: In the first half of the 19th century, steam locomotives were mainly imported to Russia from abroad. Question word: When Answer: 19th century Transformed Question: When were steam locomotives mainly imported into Russia from abroad?